A commetee is choosing a site for a community centre. The centre need to be located the same distance from the two main towns in the country. On a map , these towns have coordinates (2,14) and (16,4). Determine an equation for the line that shows all possible sites for the community centre)
Answer by CubeyThePenguin(3113)

We want to find the perpendicular bisector of the line that connects the towns.

(2, 14) and (16, 4)
slope = (4 - 14)/(16 - 2) = -10/14 = -5/7
midpoint = (2 + 16, 14 + 4) = (9, 9)

Perpendicular bisector:
slope = 7/5, passes through (9, 9)

y - 9 = 7/5 * (x - 9)